OpenStudy (anonymous):

how do i solve cos^2(theta)+sin(theta)=1?

myininaya (myininaya):

\[1-\sin^2(\theta)+\sin(\theta)-1=0\] \[-\sin^2(\theta)+\sin(\theta)=0\] \[-\sin(\theta)(\sin(\theta)-1)=0\]

OpenStudy (anonymous):

did you mean the trig identity: cos^2(theta)+sin^2(theta)=1

OpenStudy (anonymous):

probably not ...

OpenStudy (anonymous):

the theta isnt part of the exponent and its just sin(theta)

myininaya (myininaya):

\[-\sin(\theta)=0=> \sin(\theta)=0=> \theta= n \pi, n \in \mathbb{Z}\] \[\sin(\theta)-1=0=> \sin(\theta)=1=> \theta=\frac{\pi}{2}+2n \pi, \theta=\frac{3\pi}{2}+2n \pi\]

OpenStudy (anonymous):

not sin^2

myininaya (myininaya):

any questions?

OpenStudy (anonymous):

yes why did u use a negative identity ?

myininaya (myininaya):

negative?

myininaya (myininaya):

whats that mean?

myininaya (myininaya):

i mean what do you mean by negative identity

OpenStudy (anonymous):

i thought you take sin(theta) and put it on the other side

myininaya (myininaya):

\[\cos^2(\theta)=1-\sin^2(\theta)\] <- this is the only identity i used

OpenStudy (anonymous):

oh so u substituted 1-sin^2 where cos^2 is

myininaya (myininaya):

yes

OpenStudy (anonymous):

thanks can u show me the answers i would get?

myininaya (myininaya):

? i already wrote answers

OpenStudy (anonymous):

ya thanks c them now